OpenStudy (anonymous):

write the expression below as a single logarithm. show work... 2In(x+4)-In(x^2+x-12/x)

OpenStudy (anonymous):

\[\ln((x+4)^2)-\ln(\frac{x^2+x-12}{x})\] \[\ln((x+4)^2+\ln(\frac{x}{x^2+x-12})\] \[\ln(\frac{(x+4)^2\times x}{(x+4)(x-3)})\] \[\ln(\frac{x(x+4)}{x-3})\]